OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

B

OpenStudy (anonymous):

C.y = 2x – 21

OpenStudy (anonymous):

D

OpenStudy (anonymous):

@blondie16

OpenStudy (anonymous):

first solve that equation for y and tell me what you get

OpenStudy (anonymous):

thts the thing i dont know how to do it

OpenStudy (anonymous):

2x + 3y = 5 first you would sutract the 2x -2x -2x 3y=-2x+5 then you would divide by 3 _______ 3 3 3 so it would = y=-2/3x+5/3

OpenStudy (anonymous):

i get it thx i luv u (u know wat i mean)

OpenStudy (anonymous):

lol but your not done with the question! do you get the rest?

OpenStudy (anonymous):

the rest are awnsers like abcd u know

OpenStudy (anonymous):

but your not finished with the questions you have to find the perpendicular slope to the slope you got. and you have to put it through the point (9,-3)

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

so i have to plug them it in right

OpenStudy (anonymous):

lol yep :P to find the perpendicular slope you have to find the opposite reciprocal do you know how to do that?

OpenStudy (anonymous):

and no dont plug them in

OpenStudy (anonymous):

ohh thn how?

OpenStudy (anonymous):

well we found out that the slope of that line was -2/3 right? so the 'opposite' of that would be positive and the reciprocal would be 3/2 so that is the new slope: 3/2(m). now for the new point. (9(x1),-3(y1)), plug it in to the equation : y-y1=m(x-x1) the 9 goes into the x1 spot and the -3 goes into the y1 spot. dont forget, the y1 is a negative 3 and in the equation, it subtracts it so the 3 would become positive.

OpenStudy (anonymous):

ohh isn tht point slope form

OpenStudy (anonymous):

yesh it is :)

OpenStudy (anonymous):

i get it thx

OpenStudy (anonymous):

yay :D your welcome glad i could help reaper :P

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

swear i learn more here then on my school

OpenStudy (anonymous):

oops! i read perpendicular :O it is parallel :( so it would still be -2/3 sorry i misread that and lol

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

alright so if you want me to check it i will if you want haha

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome :)